OpenStudy (anonymous):

If the tangent line to y = f(x) at (6, 3) passes through the point (0, 2), find f(6) and f'(6).

OpenStudy (dumbcow):

f(6) is given as 3 f'(6) is slope of tangent line.... m = (3-2)/(6-0) = 1/6